OpenStudy (adll23):

Solve for x: 2 over 3 (x - 2) = 4x. (4 points) negative 2 over 5 negative 5 over 2 2 over 5

OpenStudy (alexandervonhumboldt2):

2/3*(x-2)=4x first multiply both sides by 3/2. we get 2/3*3/2*(x-2)=4x

OpenStudy (alexandervonhumboldt2):

woops

OpenStudy (alexandervonhumboldt2):

2/3*3/2*(x-2)=4x*3/2

OpenStudy (alexandervonhumboldt2):

now we have x-2=6x -2=5x x=-2/5

OpenStudy (agl202):

Lol, u got me a little confused there btw Welcome back!

OpenStudy (adll23):

thank you!

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (alexandervonhumboldt2):

thx @Agl202